### v1

@@ 3.1 A reduction method for NLSDP
${\mathcal{B}}d:=\begin{pmatrix}h^{\prime}(\bar{x})d;\\[2.84526pt] (P^{\top}_{\beta_{1}}[g^{\prime}_{1}(\bar{x})d]P_{\gamma_{1}};P^{\top}_{\gamma_{1}}[g^{\prime}_{1}(\bar{x})d]P_{\gamma_{1}})\\ \vdots\\ (P^{\top}_{\beta_{J}}[g^{\prime}_{J}(\bar{x})d]P_{\gamma_{J}};P^{\top}_{\gamma_{J}}[g^{\prime}_{J}(\bar{x})d]P_{\gamma_{J}})\end{pmatrix},\quad d\in{\mathcal{X}}.$ (3.10)
@@
From (2.15) one can see that ${\sf ker}({\mathcal{B}})={\sf aff}({\mathcal{C}}(\bar{x}))$, which is a finite-dimensional subspace of $\mathcal{X}$. Let $r$ be the dimension of ${\sf ker}({\mathcal{B}})$. Then, one can find a collection of linearly independent vectors $\omega_{1},\ldots,\omega_{r}\in{\mathcal{X}}$ such that ${\sf span}\{\omega_{1},\ldots,\omega_{r}\}={\sf ker}({\mathcal{B}})$. Moreover, one can define the linear operator ${\mathcal{W}}:{\mathds{R}}^{r}\to{\sf ker}({\mathcal{B}})$ via
@@
Consequently, one has ${\sf rge}({\mathcal{W}})={\sf ker}({\mathcal{B}})$. Based on the definition of ${\mathcal{W}}$ in (3.11) one can get the following result from Proposition 3.1.
@@
Let $\bar{x}$ be a nondegenerate locally optimal solution of problem (1.8) such that $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\ldots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$. For $j\in\{1,\ldots,J\}$, define $A_{j}:=g_{j}(\bar{x})+\bar{\Gamma}_{j}$ and write its eigenvalue decomposition $A_{j}=P_{j}\Lambda^{j}P_{j}^{\top}$ as in (2.4) with $P_{j}=((P_{j})_{\alpha_{j}},(P_{j})_{\beta_{j}},(P_{j})_{\gamma_{j}})$ being the corresponding orthogonal matrix and $\Lambda^{j}$ being the corresponding diagonal matrix of eigenvalues. Let ${\mathcal{Q}}$ be the linear operator defined by (2.16) and ${\mathcal{W}}$ be the linear operator defined by (3.11). Then, the mapping ${\mathbb{S}}_{\rm GE}$ defined by (3.1) has the Aubin property at 00 for $\bar{x}$ if and only if
$\begin{array}[]{ll}\displaystyle{\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}\nu\notin\Bigg{\{}\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\widetilde{U}_{\beta_{j}\beta_{j}}^{j}(P_{j})_{\beta_{j}}^{\top}\\[14.22636pt] \displaystyle\quad\big{|}\ \widetilde{U}_{\beta_{j}\beta_{j}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)((P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x}){\mathcal{W}}\nu)(P_{j})_{\beta_{j}})\Bigg{\}}\quad\forall\nu\in{\mathds{R}}^{r}\backslash\{0\}.\end{array}$ (3.12)
@@
Denote $\bar{\mu}:={\mathcal{Q}}{\mathcal{W}}\bar{\nu}$. From (3.14) one has that
@@
Recall that the linear operator ${\mathcal{B}}$ in (3.10) is well-defined. Moreover, for any $d\in{\mathcal{X}}$, $\xi\in{\mathds{R}}^{m}$, $U^{j}_{\beta_{j}\gamma_{j}}\in{\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|}$, $U^{j}_{\gamma_{j}\gamma_{j}}\in{\mathcal{S}}^{|\gamma_{j}|}$, $j=1,\ldots,J$, it holds that
@@
$\begin{array}[]{l}\displaystyle\left\langle{\mathcal{B}}^{*}\left(\xi;\big{(}U^{1}_{\beta_{j}\gamma_{j}};U^{1}_{\gamma_{j}\gamma_{j}}\big{)};\cdots;\big{(}U^{J}_{\beta_{J}\gamma_{J}};U^{J}_{\gamma_{J}\gamma_{J}}\big{)}\right),d\right\rangle\\[11.38109pt] =\left\langle\left(\xi;\big{(}U^{1}_{\beta_{j}\gamma_{j}};U^{1}_{\gamma_{j}\gamma_{j}}\big{)};\cdots;\big{(}U^{J}_{\beta_{J}\gamma_{J}};U^{J}_{\gamma_{J}\gamma_{J}}\big{)}\right),{\mathcal{B}}d\right\rangle\\[11.38109pt] \displaystyle=\langle\xi,h^{\prime}(\bar{x})d\rangle\\[2.84526pt] \qquad\displaystyle+\sum_{j=1}^{J}\langle U^{j}_{\beta_{j}\gamma_{j}},(P_{j})^{\top}_{\beta_{j}}[g_{j}^{\prime}(\bar{x})d](P_{j})_{\gamma_{j}}\rangle+\sum_{j=1}^{J}+\langle U^{j}_{\gamma_{j}\gamma_{j}},(P_{j})^{\top}_{\gamma_{j}}[g_{j}^{\prime}(\bar{x})d](P_{j})_{\gamma_{j}}\rangle\\[17.07164pt] =\langle\xi,h^{\prime}(\bar{x})d\rangle\\[2.84526pt] \displaystyle\qquad+\sum_{j=1}^{J}\left\langle(P_{j})_{\beta_{j}}U^{j}_{\beta_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}},g_{j}^{\prime}(\bar{x})d\right\rangle+\sum_{j=1}^{J}\left\langle(P_{j})_{\gamma_{j}}U^{j}_{\gamma_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}},g_{j}^{\prime}(\bar{x})d\right\rangle\\[11.38109pt] \displaystyle=\Big{\langle}\nabla h(\bar{x})\xi+\sum_{j=1}^{J}\nabla g(\bar{x})(P_{j})_{\beta_{j}}U^{j}_{\beta_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}}+\sum_{j=1}^{J}\nabla g(\bar{x})(P_{j})_{\gamma_{j}}U^{j}_{\gamma_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}},d\Big{\rangle}.\end{array}$ (3.16)
Note that ${\sf ker}({\mathcal{W}}^{*})={\sf rge}({\mathcal{B}}^{*})=({\sf ker}({\mathcal{B}}))^{\perp}$. Then by combining (3.15) and (3.16) one can get that
@@
$\begin{array}[]{ll}\displaystyle\bar{\mu}\in{\sf rge}(\nabla h(\bar{x}))-\Bigg{\{}\sum_{j=1}^{J}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}\cup\gamma_{j}}\begin{pmatrix}\widetilde{U}_{\beta_{j}\beta_{j}}^{j}&\widetilde{U}^{j}_{\beta_{j}\gamma_{j}}\\[5.69054pt] ({\widetilde{U}}^{j}_{\beta_{j}\gamma_{j}})^{\top}&\widetilde{U}^{j}_{\gamma_{j}\gamma_{j}}\end{pmatrix}(P_{j})_{\beta_{j}\cup\gamma_{j}}^{\top}\\[22.76219pt] \displaystyle\hfill\ \bigg{|}\ \begin{array}[]{ll}\widetilde{U}^{j}_{\beta_{j}\beta_{j}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)\big{(}(P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x}){\mathcal{W}}\bar{\nu})(P_{j})_{\beta_{j}}\big{)}\\[5.69054pt] \forall\,{\widetilde{U}^{j}}_{\beta_{j}\gamma_{j}}\in{\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|},{\widetilde{U}^{j}}_{\gamma_{j}\gamma_{j}}\in{\mathcal{S}}^{|\gamma_{j}|},\ j=1,\ldots,J\end{array}\Bigg{\}},\end{array}$
@@
We can readily get the following result based on Proposition 3.2.
@@
Suppose that the assumptions in Proposition 3.2 hold. Then the linear operator ${\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}$ is a nonsingular matrix in ${\mathcal{S}}^{r}$.
@@
As a result, one has ${\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}\bar{\nu}=0=-\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\widetilde{U}^{j}_{\beta_{j}\beta_{j}}(P_{j})^{\top}_{\beta_{j}}$, which contradicts (3.12). Consequently, ${\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}$ is nonsingular. ∎
@@ 3.2 Aubin property implies SSOSC
The following lemma comes from [10, Corollary 25].
Lemma 3.2.
Let $\bar{x}$ be a locally optimal solution of problem (1.8) with $\bar{y}\in{\mathcal{M}}(\bar{x})$. Suppose that the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ in (1.6) holds at $(0,0)$ for $(\bar{x},\bar{y})$. Then, $\bar{x}$ is nondegenerate, i.e., (2.14) holds, and the second-order sufficient condition (2.17) in Definition 2.1 holds with ${\mathcal{Q}}$ being defined in (2.16).
Here, we provide a remarkable improvement to Lemma 3.2.
Proposition 3.3.
Let $\bar{x}$ be a locally optimal solution to the NLSDP problem (1.8) and $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\dots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$ be a multiplier at $\bar{x}$. Suppose that the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.6) has the Aubin property at $(0,0)$ for $(\bar{x},\bar{y})$. Then, the SSOSC (2.18) in Definition 2.1 holds at $(\bar{x};\bar{y})$ with ${\mathcal{Q}}$ being defined in (2.16).
Proof.
From Lemma 3.2 we know that the constraint nondegeneracy (2.14) holds, implying that ${\mathcal{M}}(\bar{x})$ is a singleton, and that the second-order sufficient condition (2.17) holds. For all $j\in\{1,\dots,J\}$, define $A_{j}:=g_{j}(\bar{x})+\bar{\Gamma}_{j}$ and write its eigenvalue decomposition as in (2.4) with $P_{j}=((P_{j})_{\alpha_{j}},(P_{j})_{\beta_{j}},(P_{j})_{\gamma_{j}})$ being the corresponding orthogonal matrix and $\Lambda^{j}$ being the corresponding diagonal matrix of eigenvalues. Then the linear operators ${\mathcal{B}}$ in (3.10) and ${\mathcal{W}}$ in (3.11) are well-defined. Consequently, (3.12) in Proposition 3.2 holds.
Define ${\mathcal{A}}:{\mathds{R}}^{r}\to{\mathcal{S}}^{|\beta_{1}|}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}$ by ${\mathcal{A}}\nu:=({\mathcal{A}}_{1}\nu;\dots;{\mathcal{A}}_{J}\nu)$, where the linear operators ${\mathcal{A}}_{j}:{\mathds{R}}^{r}\to{\mathcal{S}}^{|\beta_{j}|}$, $j=1,\ldots,J$, are defined by
${\mathcal{A}}_{j}\nu:=(P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x}){\mathcal{W}}\nu)(P_{j})_{\beta_{j}},\quad\nu\in{\mathds{R}}^{r}.$ (3.17)
@@
Since (2.14) holds, for any given $Y_{j}\in{\mathcal{S}}^{p_{j}},\,j=1,\dots,J$, one can always find a vector $d\in{\mathcal{X}}$ and the matrices
@@
Then by (2.15) we know that $d\in{\sf aff}(C(\bar{x}))$. Recall that ${\sf rge}({\mathcal{W}})={\sf ker}({\mathcal{B}})={\sf aff}({\mathcal{C}}(\bar{x}))$. Then there exists a vector $\nu\in{\mathds{R}}^{r}$ such that $d={\mathcal{W}}\nu$. Since $(Y_{j})_{\beta_{j}\beta_{j}}$ can be any matrix in ${\mathcal{S}}^{|\beta_{j}|}$, one can see that the linear operator ${\mathcal{A}}$ defined by (3.17) is surjective. Recall from (2.11) and (2.12) that
${\mathcal{C}}(\bar{x})=\left\{d\in{\mathcal{X}}\ \Bigg{|}\ \begin{array}[]{ll}h^{\prime}(\bar{x})d=0,\,(P_{j})^{\top}_{\beta_{j}}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\beta_{j}}\succeq 0,\\[2.84526pt] (P_{j})^{\top}_{\beta_{j}}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\gamma_{j}}=0,\\[2.84526pt] (P_{j})^{\top}_{\gamma_{j}}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\gamma_{j}}=0,\,\,j=1,\dots,J\end{array}\right\}\subseteq{\sf rge}({\mathcal{W}}).$
Then the second-order sufficient condition (2.17) can be equivalently recast as
Also, the SSOSC (2.18) can be equivalently written as
Moreover, since ${\mathcal{A}}$ is surjective, one can get from Lemma 2.2 that
$\Omega^{\circ}$ =$=\{{\mathcal{A}}^{*}(\Theta_{1};\dots;\Theta_{J})\mid\Theta_{j}\in-{\mathcal{S}}^{|\beta_{j}|}_{+},\ \,j=1,\dots,J\}$
@@
In the following, we shall show that (3.20) holds.
Suppose on the contrary that (3.20) does not hold. Consider the following auxiliary optimization problem

### v2

@@ 3.1 A reduction method for NLSDP
${\mathcal{B}}d:=\begin{pmatrix}h^{\prime}(\bar{x})d\\[2.84526pt] \big{(}{2}P^{\top}_{\beta_{1}}[g^{\prime}_{1}(\bar{x})d]P_{\gamma_{1}};P^{\top}_{\gamma_{1}}[g^{\prime}_{1}(\bar{x})d]P_{\gamma_{1}}\big{)}\\[2.84526pt] \vdots\\[2.84526pt] \big{(}{2}P^{\top}_{\beta_{J}}[g^{\prime}_{J}(\bar{x})d]P_{\gamma_{J}};P^{\top}_{\gamma_{J}}[g^{\prime}_{J}(\bar{x})d]P_{\gamma_{J}}\big{)}\end{pmatrix},\quad d\in{\mathcal{X}}.$ (3.11)
@@
From (2.15) one can see that ${\sf ker}({\mathcal{B}})={\sf aff}({\mathcal{C}}(\bar{x}))$, which is a finite-dimensional subspace of $\mathcal{X}$. Let $r$ be the dimension of ${\sf ker}({\mathcal{B}})$. Then, a collection of linearly independent vectors $\omega_{1},\ldots,\omega_{r}\in{\mathcal{X}}$ can be found such that ${\sf span}\{\omega_{1},\ldots,\omega_{r}\}={\sf ker}({\mathcal{B}})$. Moreover, one can define the linear operator ${\mathcal{W}}:{\mathds{R}}^{r}\to{\sf ker}({\mathcal{B}})$ via
@@
Consequently, ${\sf rge}({\mathcal{W}})={\sf ker}({\mathcal{B}})={\sf aff}({\mathcal{C}}(\bar{x}))$. Based on the definition of ${\mathcal{W}}$ in (3.12) we have the following result from Proposition 3.1.
@@
Suppose that the conditions of Proposition 3.1 are satisfied. Let ${\mathcal{Q}}$ be the linear operator defined by (2.16) and ${\mathcal{W}}$ be the linear operator defined by (3.12). Then, the mapping ${\mathbb{S}}_{\rm GE}$ defined by (3.1) has the Aubin property at 00 for $\bar{x}$ if and only if
$\begin{array}[]{ll}\displaystyle{\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}\nu\notin\Bigg{\{}{-}\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\widetilde{U}_{\beta_{j}\beta_{j}}^{j}(P_{j})_{\beta_{j}}^{\top}\\[14.22636pt] \displaystyle\qquad\big{|}\ \widetilde{U}_{\beta_{j}\beta_{j}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)((P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x}){\mathcal{W}}\nu)(P_{j})_{\beta_{j}})\Bigg{\}}\quad\forall\nu\in{\mathds{R}}^{r}\backslash\{0\}.\end{array}$ (3.13)
@@
Denote $\bar{\mu}:={\mathcal{Q}}{\mathcal{W}}\bar{\nu}$. From (3.15) one has
@@
Recall that the linear operator ${\mathcal{B}}$ in (3.11) is well-defined and
${\sf ker}({\mathcal{W}}^{*})={({\sf rge}({\mathcal{W}}))^{\perp}=}({\sf ker}({\mathcal{B}}))^{\perp}={\sf rge}({\mathcal{B}}^{*}).$
@@
Thus, we calculate ${\sf rge}({\mathcal{B}}^{*})$ to further reformulate (3.16). Note that for any $d\in{\mathcal{X}}$, $\xi\in{\mathds{R}}^{m}$, $U^{j}_{\beta_{j}\gamma_{j}}\in{\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|}$, $U^{j}_{\gamma_{j}\gamma_{j}}\in{\mathcal{S}}^{|\gamma_{j}|}$, $j=1,\ldots,J$, it holds from (3.11) that
@@
$\begin{array}[]{l}\displaystyle\left\langle{\mathcal{B}}^{*}\left(\xi;\big{(}U^{1}_{\beta_{j}\gamma_{j}};U^{1}_{\gamma_{j}\gamma_{j}}\big{)};\cdots;\big{(}U^{J}_{\beta_{J}\gamma_{J}};U^{J}_{\gamma_{J}\gamma_{J}}\big{)}\right),d\right\rangle\\[5.69054pt] =\left\langle\left(\xi;\big{(}U^{1}_{\beta_{j}\gamma_{j}};U^{1}_{\gamma_{j}\gamma_{j}}\big{)};\cdots;\big{(}U^{J}_{\beta_{J}\gamma_{J}};U^{J}_{\gamma_{J}\gamma_{J}}\big{)}\right),{\mathcal{B}}d\right\rangle\\[11.38109pt] \displaystyle=\langle\xi,h^{\prime}(\bar{x})d\rangle\\ \qquad\displaystyle+\sum_{j=1}^{J}\left\langle U^{j}_{\beta_{j}\gamma_{j}},2(P_{j})^{\top}_{\beta_{j}}[g_{j}^{\prime}(\bar{x})d](P_{j})_{\gamma_{j}}\right\rangle+\sum_{j=1}^{J}\left\langle U^{j}_{\gamma_{j}\gamma_{j}},(P_{j})^{\top}_{\gamma_{j}}[g_{j}^{\prime}(\bar{x})d](P_{j})_{\gamma_{j}}\right\rangle\\[17.07164pt] \displaystyle=\langle\nabla h(\bar{x})\xi,d\rangle\\ \qquad\displaystyle+2\sum_{j=1}^{J}\left\langle(P_{j})_{\beta_{j}}U^{j}_{\beta_{j}\gamma_{j}}(P_{j})_{\gamma_{j}}^{\top},g_{j}^{\prime}(\bar{x})d\right\rangle+\sum_{j=1}^{J}\left\langle(P_{j})_{\gamma_{j}}U^{j}_{\gamma_{j}\gamma_{j}}(P_{j})_{\gamma_{j}}^{\top},g_{j}^{\prime}(\bar{x})d\right\rangle.\end{array}$
We know that $g_{j}^{\prime}(\bar{x})d$ is symmetric for all $j=1,\ldots J$. Therefore, one has
$\begin{array}[]{l}\displaystyle\left\langle{\mathcal{B}}^{*}\left(\xi;\big{(}U^{1}_{\beta_{j}\gamma_{j}};U^{1}_{\gamma_{j}\gamma_{j}}\big{)};\cdots;\big{(}U^{J}_{\beta_{J}\gamma_{J}};U^{J}_{\gamma_{J}\gamma_{J}}\big{)}\right),d\right\rangle\\[11.38109pt] \displaystyle=\langle\nabla h(\bar{x})\xi,d\rangle+\sum_{j=1}^{J}\left\langle(P_{j})_{\gamma_{j}}(U^{j}_{\beta_{j}\gamma_{j}})^{\top}(P_{j})^{\top}_{\beta_{j}},g_{j}^{\prime}(\bar{x})d\right\rangle\\ \displaystyle\qquad+\sum_{j=1}^{J}\left\langle(P_{j})_{\beta_{j}}U^{j}_{\beta_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}},g_{j}^{\prime}(\bar{x})d\right\rangle+\sum_{j=1}^{J}\left\langle(P_{j})_{\gamma_{j}}U^{j}_{\gamma_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}},g_{j}^{\prime}(\bar{x})d\right\rangle\\ \displaystyle=\Big{\langle}\nabla h(\bar{x})\xi+\sum_{j=1}^{J}\nabla g_{j}(\bar{x})(P_{j})_{\gamma_{j}}(U^{j}_{\beta_{j}\gamma_{j}})^{\top}(P_{j})^{\top}_{\beta_{j}}\\ \qquad\displaystyle+\sum_{j=1}^{J}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}U^{j}_{\beta_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}}+\sum_{j=1}^{J}\nabla g_{j}(\bar{x})(P_{j})_{\gamma_{j}}U^{j}_{\gamma_{j}\gamma_{j}}(P_{j})^{\top}_{\gamma_{j}},d\Big{\rangle}.\end{array}$ (3.17)
Then by combining (3.16) and the formula of ${\sf rge}({\mathcal{B}}^{*})\equiv{\sf ker}({\mathcal{W}}^{*})$ given by (3.17) one can get
@@
$\begin{array}[]{ll}\displaystyle\bar{\mu}\in{\sf rge}\big{(}\nabla h(\bar{x})\big{)}-\Bigg{\{}\sum_{j=1}^{J}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}\cup\gamma_{j}}\begin{pmatrix}\widetilde{U}_{\beta_{j}\beta_{j}}^{j}&\widetilde{U}^{j}_{\beta_{j}\gamma_{j}}\\[5.69054pt] ({\widetilde{U}}^{j}_{\beta_{j}\gamma_{j}})^{\top}&\widetilde{U}^{j}_{\gamma_{j}\gamma_{j}}\end{pmatrix}(P_{j})_{\beta_{j}\cup\gamma_{j}}^{\top}\\[22.76219pt] \displaystyle\hfill\ \bigg{|}\ \begin{array}[]{ll}\widetilde{U}^{j}_{\beta_{j}\beta_{j}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)\big{(}(P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x}){\mathcal{W}}\bar{\nu})(P_{j})_{\beta_{j}}\big{)}\\[5.69054pt] \forall\,{\widetilde{U}^{j}}_{\beta_{j}\gamma_{j}}\in{\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|},{\widetilde{U}^{j}}_{\gamma_{j}\gamma_{j}}\in{\mathcal{S}}^{|\gamma_{j}|},\ j=1,\ldots,J\end{array}\Bigg{\}},\end{array}$
@@
Propositions 3.1 and 3.2 constitute the reduction method by reformulating the Mordukhovich criterion to characterize the Aubin property of ${\mathbb{S}}_{\rm GE}$ defined by (3.1) to that given by (3.13), in which only the indices $\beta_{j}$ are involved while the indices $\alpha_{j}$ and $\gamma_{j}$ are eliminated. Such a reduction provides a more accessible form of the Aubin property, which is convenient for the discussions on obtaining the SSOSC from the second-order sufficient condition. Before that, we provide a useful result based on Proposition 3.2.
@@
Under the conditions of Proposition 3.2, the linear operator ${\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}$ is a nonsingular matrix in ${\mathcal{S}}^{r}$.
@@
As a result, one has ${\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}\bar{\nu}=0=-\sum_{j=1}^{J}{\mathcal{W}}^{*}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}}\widetilde{U}^{j}_{\beta_{j}\beta_{j}}(P_{j})^{\top}_{\beta_{j}}$, which contradicts (3.13). Consequently, ${\mathcal{W}}^{*}{\mathcal{Q}}{\mathcal{W}}$ is not singular. ∎
@@ 3.2 Aubin property implies SSOSC
Note that Proposition 3.2 in the previous subsection provides a characterization of the Aubin property of the mapping ${\mathbb{S}}_{\rm GE}$ defined by (3.1) using only the indices $\beta_{j}$. In this part, our objective is to utilize (3.13) and the second-order sufficient condition in (2.17) (as a consequence of the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ for locally optimal solutions by Lemma 2.2) to derive the SSOSC in (2.18). For this purpose, we need to first reformulate the two second-order optimality conditions to pave the way for using Proposition 3.2.
Under the conditions of Proposition 3.2, define a linear operator ${\mathcal{A}}:{\mathds{R}}^{r}\to{\mathcal{S}}^{|\beta_{1}|}\times\dots\times{\mathcal{S}}^{|\beta_{J}|}$ by
${\mathcal{A}}\nu:=({\mathcal{A}}_{1}\nu;\dots;{\mathcal{A}}_{J}\nu),$ (3.18)
where the linear operators ${\mathcal{A}}_{j}:{\mathds{R}}^{r}\to{\mathcal{S}}^{|\beta_{j}|}$, $j=1,\ldots,J$, are given by
${\mathcal{A}}_{j}\nu:=(P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x}){\mathcal{W}}\nu)(P_{j})_{\beta_{j}},\quad\nu\in{\mathds{R}}^{r}.$
@@
Note that $\bar{x}$ is nondegenerate, and the linear operators ${\mathcal{B}}$ in (3.11) and ${\mathcal{W}}$ in (3.12) are well defined with ${\sf rge}({\mathcal{W}})={\sf ker}({\mathcal{B}})={\sf aff}({\mathcal{C}}(\bar{x}))$. Thus, for any $d\in{\sf aff}(C(\bar{x}))$ there exists a vector $\nu\in{\mathds{R}}^{r}$ such that $d={\mathcal{W}}\nu$. Consequently, the SSOSC (2.18) can be equivalently written as
Moreover, recall from (2.11), (2.12) and (2.15) that
$\begin{array}[]{ll}{\mathcal{C}}(\bar{x})&=\left\{d\in{\mathcal{X}}\ \Bigg{|}\ \begin{array}[]{ll}h^{\prime}(\bar{x})d=0,\,(P_{j})^{\top}_{\beta_{j}}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\beta_{j}}\succeq 0,\\[2.84526pt] (P_{j})^{\top}_{\beta_{j}}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\gamma_{j}}=0,\\[2.84526pt] (P_{j})^{\top}_{\gamma_{j}}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\gamma_{j}}=0,\,\,j=1,\dots,J\end{array}\right\}\\[17.07164pt] &=\left\{d\in{\mathcal{X}}\ \big{|}\ d\in{\sf aff}({\mathcal{C}}(\bar{x})),\,(P_{j})^{\top}_{\beta_{j}}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\beta_{j}}\succeq 0,j=1,\dots,J\right\}.\end{array}$
Therefore, the second-order sufficient condition (2.17) can be equivalently recast as
In the following, we show that the conditions of Lemma 2.2 can actually imply (3.19), hence the SSOSC given by (2.18), which constitutes a remarkable improvement to the second-order sufficient condition in this Lemma. To achieve this, we need the explicit formula of $\Omega^{\circ}$, and the following result is essential.
Lemma 3.2.
Under the conditions of Proposition 3.2, the linear operator ${\mathcal{A}}$ defined by (3.18) is surjective.
Proof.
Since the constraint nondegeneracy (2.14) holds, for any given $Y_{j}\in{\mathcal{S}}^{p_{j}},\,j=1,\dots,J$, one can always find a vector $d\in{\mathcal{X}}$ and the matrices
@@
Since $(Y_{j})_{\beta_{j}\beta_{j}}$ can be any matrix in ${\mathcal{S}}^{|\beta_{j}|}$, one can see that the linear operator ${\mathcal{A}}$ defined by (3.18) is subjective. This completes the proof. ∎
Since ${\mathcal{A}}$ is subjective by Lemma 3.2, one can explicitly calculate from Lemma 2.3 that
$\Omega^{\circ}$ =$=\{{\mathcal{A}}^{*}(\Theta_{1};\dots;\Theta_{J})\mid\Theta_{j}\in-{\mathcal{S}}^{|\beta_{j}|}_{+},\ \,j=1,\dots,J\}$ (3.22)
@@
Now, we are ready to present the main result of this section.
Theorem 3.1.
Let $\bar{x}$ be a locally optimal solution to the NLSDP problem (1.8) and $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\dots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$ be a multiplier at $\bar{x}$. Suppose that the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.6) has the Aubin property at $(0;0)$ for $(\bar{x};\bar{y})$. Then, the SSOSC (2.18) in Definition 2.1 holds at $(\bar{x};\bar{y})$ with ${\mathcal{Q}}$ being defined in (2.16).
Proof.
Following the notation of Proposition 3.2, for $j\in\{1,\ldots,J\}$, define $A_{j}:=g_{j}(\bar{x})+\bar{\Gamma}_{j}$ and write its eigenvalue decomposition $A_{j}=P_{j}\Lambda^{j}P_{j}^{\top}$ as in (2.4) with $P_{j}=\big{(}(P_{j})_{\alpha_{j}},(P_{j})_{\beta_{j}},(P_{j})_{\gamma_{j}}\big{)}$ being the corresponding orthogonal matrix and $\Lambda^{j}$ being the corresponding diagonal matrix of eigenvalues. Then, from the above analysis, we only need to prove that (3.21) holds.
We start by considering the following auxiliary optimization problem